A weight of 20kg acts tangentially at a radius of 650 mm and causes a pulley to accelerate uniformly from 35 rev/min to 120 rev/min in 20 seconds. Calculate the work done and the average power required to cause this acceleration.


Clearly show the following values

wi=?

wf=?

θ=?

Torque=?

W=?

Pave=?



1
Expert's answer
2021-09-25T11:38:59-0400

"W=20kg"

t=20 sec


actual air temprature Ta ="25\degree"

from the data

"h\alpha 0.8 V"

"h=k(0.8V)"

in the thermocouple read the radiation from the wall is proportional to "\Delta T"

"q_{radiation}"

"q_{radiation} = C(\Delta T)"

=C(Tb-Tw)

consider energy balance

"q_{conduction}"

"h\Delta T=C \Delta (Tb -Tw)"

"k(0.8V)\Delta T = C (Tb - Tw)"

from equation 1

"0.8 (5) k (25 - 50) = C (50 - Tw)"

"0.8 (10) k (25 - 40) = C (40 - Tw)"

"2\/3 =(0.8) (5) k (-25) \/(0.8)(10) k(15) = C (50 - Tw) \/ C ( 40 - Tw )"